. Assume that the demand and supply functions for a commodity is given as P = 80 - Q and P = 20 + 2Q , respectively. If the seller must pay a tax of $6 per unit. Calculate the new supply function, equilibrium price (with tax and without tax) and the equilibrium quantity.


1
Expert's answer
2022-02-07T17:22:27-0500

1) Rewritten Supply Function

Supply function before tax considerations => P = 20 + 2Q

Tax per unit of output = $6

Supply function after tax => P = (20 + 2Q) + 6

=> P = 26 + 2Q


2) Equilibrium Quantity (Without Tax)

80 - Q = 20 + 2Q

80 - 20 = 2Q + Q

60 = 3Q

Q = 20

Equilibrium Quantity (With Tax)

80 - Q = 26 + 2Q

80 - 26 = 2Q + Q

54 = 3Q

Q = 18


3) Equilibrium Price (Without Tax)

P = 80 - Q

If Q = 20 at equilibrium,

P = 80 - 20

P = 60

Equilibrium Price (With Tax)

P = 80 - Q

If Q = 18 at equilibrium,

P = 80 - 18

P = 62